OpenStudy (lena772):

The sales totals at Macy's food store have increased exponentially over the month. Which of these best shows the sales in the first three months? $1200 in the first month, $1220 in the second month, $1240 in the third month $1200 in the first month, $1272 in the second month, $1344 in the third month $1200 in the first month, $1285 in the second month, $1370.96 in the third month $1200 in the first month, $1224 in the second month, and $1248.48 in the third month

OpenStudy (kittiwitti1):

1. Is this multiple choice? 2. Where is the diagram? There should be a diagram to go along with this. The above is not sufficient information.

OpenStudy (lena772):

It is multiple choice but there is no diagram

OpenStudy (kittiwitti1):

Hmmm. According to my calculator the exponents for any one of them would be a decimal...

OpenStudy (kittiwitti1):

Let me try again on paper.

OpenStudy (lena772):

Okay :)

OpenStudy (agent0smith):

Notice that in two of them, the $ amount increases the same each time... that's not exponential growth. So eliminate those 2.

OpenStudy (kittiwitti1):

So the formula should be \[log_ax=y\]with y being the value in the second month... and x being 1200....We should find a with those known values.

OpenStudy (lena772):

not A or B right @agent0smith

OpenStudy (agent0smith):

Correct, both linear.

OpenStudy (kittiwitti1):

Sorry, wrong formula.\[a^x=y,~a^{1200}=y\]

OpenStudy (lena772):

I'm leaning towards C

OpenStudy (agent0smith):

If exponential, there should be a common multiple between the numbers... divide one by the number before it to check there is $1200 in the first month, $1285 in the second month, $1370.96 in the third month for this one, 1285/1200 = 1.0708, but 1370.96 /1285 = 1.06689

OpenStudy (agent0smith):

the last one, 1224/1200 = 1.02, and 1248.48/1224 = 1.02 Which do you think is exponential?

OpenStudy (kittiwitti1):

Throw the guy a medal for being awesome. :D I could never do that.

OpenStudy (lena772):

The last one of course.

OpenStudy (agent0smith):

Remember exponential means something like a^n... means you're multiplying by a over and over.

OpenStudy (lena772):

Exponential is a ratio

OpenStudy (agent0smith):

GJ.

OpenStudy (lena772):

Thank you so much!

OpenStudy (kittiwitti1):

All credit goes to @agent0smith I barely did anything

OpenStudy (anonymous):

So is this D?

OpenStudy (lena772):

yes.